# A primer on inference and prediction with epidemic renewal models and sequential Monte Carlo

### v1

@@ 6.3 Model 3: Reporting biases, temporal effects, and projections
To account for reporting delays, the expected number of cases $\mu_{t}$ at time $t$ is modelled as a function of past incidence $I_{1:t-1}$ and an incubation period PMF $d_{u}$, assumed to be Gamma-distributed with mean 5.5 days and standard deviation 2.3 days (an incubation period distribution previously used for COVID-19 modelling in New Zealand64, also discretised by evaluating at $t=1,2,\ldots$):
@@
Using PMMH (algorithm 2), we find that all three models produce similar estimates of $\sigma$. The day-of-week and temporally aggregated models produce comparable estimates of $\phi$, while the estimated overdispersion in the naive model is much higher. This is expected, as the naive model attributes day-of-week noise to overdispersion in the observation process, while the other models explicitly account for this. Estimates of the day-of-week relative reporting rates from the day-of-week model are given in figure 8-E, highlighting statistically significantly greater reporting rates on weekdays (particularly Monday) than on weekends.
@@
Four-week projections are produced by iterating the estimated state-space model forward and sampling from the observation model. In all three models, 100% of future cases fall inside the 95% posterior predictive credible intervals, suggesting that all models overestimate the level of uncertainty (producing credible intervals that are too wide), most likely a result of model misspecification. Allowing for more complex correlation structures in the dynamic model for $R_{t}$ may reduce this uncertainty (for example, the assumption that $R_{t}$ follows a random walk ignores potential mean-reverting behaviour) 6.
@@ 7 Discussion
We have introduced a general framework for fitting epidemic renewal models using SMC methods and demonstrated their application in estimating hidden states, parameters, and making projections. The primary strength of these methods lies in their flexibility: they can fit any state-space model, a form that encompasses many popular epidemiological models. This allows for rapid fitting, testing, and comparison of a wide range of models, facilitating the rapid development of solutions to emerging epidemiological challenges. All code and data to reproduce all results in this paper (and more) are provided online.
We demonstrate our methods on three examples. Each example consists of a distinct model structure, dataset, and aim, yet the same general framework is used to fit each model. The first example demonstrates the estimation of $R_{t}$ using a simple renewal model, similar to EpiFilter 5, with the addition of joint estimation of the peak $R_{t}$ value and the timing of this peak. The second example demonstrates the estimation of $R_{t}$ while accounting for both reporting noise 65, 66 and imported cases 67, 68, 69, while simultaneously estimating the probability of elimination 9, 34, thus unifying the cited works. The final example demonstrates the estimation of $R_{t}$ in the presence of reporting delays 70, 15, 66 and day-of-week effects 15, 71 or temporally aggregated data 47, 72, 73, as well as the ability to produce short-term projections 8, and the importance of model comparison and selection 6, 18, again unifying the cited (and many non-cited) works. Many other extensions are possible, particularly the ability to incorporate multiple data sources 74, 33.
@@
Correctly accounting for uncertainty in model parameters has previously been shown to be critical for robust uncertainty quantification, and thus for robust policymaking 75, 40. Our focus on marginalising out parameter uncertainty is a key strength of these methods, and is not included in many other epidemiological methods, including many SMC-based approaches 32, 34.
@@
Epidemic models range from highly mechanistic (such as compartmental SIR-type models) to purely statistical (such as time series regression or exponential smoothers). The renewal model is semi-mechanistic in that it imposes a simple structure but does not model the entire underlying process. This leads to a model that is flexible, interpretable, and can produce well-calibrated short-term projections. However, like any mechanistic model, the renewal model and dynamic model on $R_{t}$ jointly imply a specific autocorrelation structure in the data. If this structure is misspecified and the only goal is the projection of observable data, then model may underperform compared to more flexible statistical models 6.
@@
Our focus on simplicity and flexibility results in algorithms that are known to be less computationally efficient than more specialised alternatives. Areas for additional consideration include the use of a better-tuned proposal distribution (in algorithm 1) 52, the use of more advanced resampling schemes (in algorithm 1) 51, and the use of better-tuned PMMH proposal densities (in algorithm 2) 43, 76. Many solutions to these problems exist within the literature, often leveraging specific details about the chosen model, and we encourage the reader to seek out more efficient algorithms once a suitable model has been specified. In addition to epidemiology-specific literature 35, 36, we also recommend two texts from the broader SMC literature 19, 51 and a tutorial paper by Doucet & Johansen (2011)77.
@@
An effective class of alternatives to SMC-type methods for performing inference and generating projections with epidemic renewal models are probabilistic-programming-language-based Gaussian process approximations, such as EpiNow2 4. These methods offer a similar level of flexibility and, in some cases, can be faster than our approach. However, these methods are more complex mathematically, whereas ours are more accessible to those without a strong mathematical background and can be implemented in a few lines of code, making them easier to debug. Furthermore, our methods can easily be adapted for online updates, allowing for real-time analysis of epidemic data, whereas Gaussian-process-based models are typically offline methods. Finally, popular probabilistic programming languages, such as Stan, do not support discrete parameters. Many quantities of interest, such as infection incidence, are discrete, which our methods can handle natively, whereas other approaches require additional approximations.
@@
Numerous methods for estimating $R_{t}$ and conducting inference on epidemic data exist, often tailored to specific pathogens or datasets. More general models have also been created, although these can be challenging to implement or are confined to pre-built software packages. While these existing packages simplify the process, they can obscure the underlying structure of the model and associated assumptions. This primer aims to strike a balance: enabling researchers to quickly and easily construct their own models from scratch, ensuring they understand the assumptions and structure of the model, while facilitating rapid testing and comparison of different models. By providing a general framework for fitting epidemic renewal models, we hope these methods contribute to the development of robust and useful epidemic models that can inform public health decision-making in future outbreaks across a range of pathogens.

### v2

@@ 6.3 Model 3: Reporting biases, temporal effects, and projections
To account for reporting delays, the expected number of cases $\mu_{t}$ at time $t$ is modelled as a function of past incidence $I_{1:t-1}$ and an incubation period PMF $d_{u}$, assumed to be Gamma-distributed with mean 5.5 days and standard deviation 2.3 days (an incubation period distribution previously used for COVID-19 modelling in New Zealand76, also discretised by evaluating at $t=1,2,\ldots$):
@@
Using PMMH (Algorithm 2), we find that all three models produce similar estimates of $\sigma$ (Table 1). The day-of-week and temporally aggregated models produce comparable estimates of $\phi$, while the estimated overdispersion in the naive model is much higher. This is expected, as the naive model attributes day-of-week noise to overdispersion in the observation process, while the other models explicitly account for this. Estimates of the day-of-week relative reporting rates from the day-of-week model are given in Figure 8-E, highlighting statistically significantly greater reporting rates on weekdays (particularly Monday) than on weekends.
@@
Four-week projections are produced by iterating the estimated state-space model forward and sampling from the observation model. In all three models, 100% of future cases fall inside the 95% posterior predictive credible intervals, suggesting that all models overestimate the level of uncertainty (producing credible intervals that are too wide), most likely a result of model misspecification. Allowing for more complex correlation structures in the dynamic model for $R_{t}$ may reduce this uncertainty (for example, the assumption that $R_{t}$ follows a random walk ignores potential mean-reverting behaviour) 8.
@@ 7 Discussion
We have introduced a general framework for fitting discrete-time epidemic renewal models using SMC methods and demonstrated their application in estimating hidden states, parameters, and making projections. The primary strength of these methods lies in their flexibility: they can fit any state-space model, a form that encompasses many popular epidemiological models. This allows for rapid fitting, testing, and comparison of a wide range of models, facilitating the rapid development of solutions to emerging epidemiological challenges. All code and data to reproduce all results in this paper (and more) are provided online.
We demonstrate our methods on three examples. Each example consists of a distinct model structure, dataset, and aim, yet the same general framework is used to fit each model. The first example demonstrates the estimation of $R_{t}$ using a simple renewal model, similar to EpiFilter 7, with the addition of joint estimation of the peak $R_{t}$ value and the timing of this peak. The second example demonstrates the estimation of $R_{t}$ while accounting for both reporting noise 77, 78 and imported cases 79, 80, 81, while simultaneously estimating the probability of elimination 11, 36, thus unifying the cited works. The final example demonstrates the estimation of $R_{t}$ in the presence of reporting delays 82, 17, 78 and day-of-week effects 17, 83 or temporally aggregated data 49, 55, 57, as well as the ability to produce short-term projections 10, and the importance of model comparison and selection 8, 20, again unifying the cited (and many non-cited) works. Many other extensions are possible, particularly the ability to incorporate multiple data sources 84, 23.
@@
Correctly accounting for uncertainty in model parameters has previously been shown to be critical for robust uncertainty quantification, and thus for robust policymaking 85, 42. Our focus on marginalising out parameter uncertainty is a key strength of these methods, and is not included in many other epidemiological methods, including many SMC-based approaches 35, 36. By coupling existing particle filter-only approaches with a PMMH algorithm like that presented here, an additional source of uncertainty (uncertainty in fixed parameters) can be accounted for, increasing the robustness of these methods.
@@
Epidemic models range from highly mechanistic (such as compartmental SEIR-type models) to purely statistical (such as time series regression or exponential smoothers). The renewal model is semi-mechanistic in that it imposes a simple structure but does not model the entire underlying process. This leads to a model that is flexible, interpretable, and can produce well-calibrated short-term projections. However, like any mechanistic model, the renewal model and dynamic model on $R_{t}$ jointly imply a specific autocorrelation structure in the data. If this structure is misspecified and the only goal is the projection of observable data, then model may underperform compared to more flexible statistical models 8.
@@
Our focus on simplicity and flexibility results in algorithms that are known to be less computationally efficient than more specialised alternatives. Areas for additional consideration include the use of a better-tuned proposal distribution (in Algorithm 1) 63, the use of more advanced resampling schemes (in Algorithm 1) 59, and the use of better-tuned PMMH proposal densities (in Algorithm 2) 45, 66. Many solutions to these problems exist within the literature, often leveraging specific details about the chosen model, and we encourage the reader to seek out more efficient algorithms once a suitable model has been specified. In addition to epidemiology-specific literature 37, 38, we also recommend two texts from the broader SMC literature 21, 59 and a tutorial paper by Doucet & Johansen (2011)62.
@@
An effective class of alternatives to SMC-type methods for performing inference and generating projections with epidemic renewal models are probabilistic-programming-language-based Gaussian process approximations, such as EpiNow2 6. These methods offer a similar level of flexibility and, in some cases, can be faster than our approach. However, these methods are more complex mathematically, whereas ours are more accessible to those without a strong mathematical background and can be implemented in a few lines of code, making them easier to debug. Furthermore, our methods can easily be adapted for online updates, allowing for real-time analysis of epidemic data, whereas Gaussian-process-based models are typically offline methods. Finally, popular probabilistic programming languages, such as Stan, do not support discrete parameters. Many quantities of interest, such as infection incidence, are discrete, which our methods can handle natively, whereas other approaches require additional approximations.
@@
Numerous methods for estimating $R_{t}$ and conducting inference on epidemic data exist, often tailored to specific pathogens or datasets. More general models have also been created, although these can be challenging to implement or are confined to pre-built software packages. While these existing packages simplify the process, they can obscure the underlying structure of the model and associated assumptions. This primer aims to strike a balance: enabling researchers to quickly and easily construct their own models from scratch, ensuring they understand the assumptions and structure of the model, while facilitating rapid testing and comparison of different models. By providing a general framework for fitting discrete-time epidemic renewal models, we hope these methods contribute to the development of robust and useful epidemic models that can inform public health decision-making in future outbreaks across a range of pathogens.